# Long-Term Fairness Inquiries and Pursuits in Machine Learning: A Survey of Notions, Methods, and Challenges

### v1

@@ Discussion
Compared with the standard static fairness interventions, the summarized mitigation methods mainly contribute to formulating the dynamics of human-model interactions and establishing the corresponding solutions based on the group utility shifts. Long-term fairness under unknown dynamics and/or without leveraging demographic information remains largely unexplored but promising, especially considering privacy laws (Chen et al. 2019) that limit access to such data. From the perspective of temporal learning, it is positive that long-term fairness is focused on the multi-step sequential learning setting. However, the current research typically only focuses on optimizing the model learner at each time step. Understanding the tensions between ensuring fairness at each time step and cumulatively at the final step needs more investigation. Finally, while RL methods are flexible and learn the dynamics of the environment well, they are data-intensive (Dulac-Arnold et al. 2021; Henderson et al. 2018) and often unsuitable for real-world scenarios. Hence, developing techniques that can efficiently estimate the dynamics is crucial. Notably, the recent advance in performative prediction (Perdomo et al. 2020) provides a new perspective on off-policy learning for modeling the interplay between the model policy and agents.
@@ 5 Evaluation of Long-Term Fairness
Application-Specific Studies
@@
Labor Market. The seminal work of (Hu and Chen 2018) considers a two-stage dynamic reputational labor market model to capture the long-term and reinforcing nature of disparate outcomes. They propose imposing demographic parity within a temporary labor market to achieve long-term equilibrium in the permanent reputational labor market based on affirmative action. Separately, Celis et al. (2021) investigate the long-term impact of the Rooney Rule (Collins 2007), a hiring policy aimed at mitigating implicit bias by mandating decision-makers to include at least one candidate from an underrepresented group. They employ a mathematical model of implicit bias, featuring two candidate groups, one of which is underrepresented. Each candidate possesses both a true, latent utility and an observed utility, subject to bias by a multiplicative factor. Candidates are then shortlisted by a selection panel based on observed utility. Their findings demonstrate that the rule progressively enhances minority hiring while maintaining overall utility.
Predictive Policing Ensign et al. (2018) explore resource allocation in predictive policing (Mohler et al. 2015), which deploys patrol officers based on historical crime data. They identify a feedback phenomenon present in similar algorithms across contexts like recidivism prediction and lending, where prediction outcomes heavily influence algorithm feedback. Using the Pólya urn model, they explain this feedback occurrence and propose black-box solutions, such as reweighing input incidents. Elzayn et al. (2019), go one step further by proposing an algorithm which, under simplification assumptions, converges to an optimally fair resource allocation, even in the absence of known stationary ground-truth crime rates, avoiding the runaway feedback loop effect.
Long-Term Fairness in Recommender Systems
Long-term fairness considerations naturally align with Recommender Systems due to evolving user interactions and item popularity, impacting system performance (Liu et al. 2021). Consequently, the RL framework is favored for its ability to model interactive dynamics. The works mostly focus on 1) user-side fairness, addressing equitable treatment for users, and 2) item-side fairness, ensuring unbiased item recommendation stemming from popularity bias due to disproportionate exposure (Chen et al. 2023).
@@
Long-term fairness efforts in recommendation systems primarily target item-side fairness, addressing the challenge by modeling the sequential interactions between consumers and recommender systems using an MDP framework. A few works in this direction adopt an Actor-Critic RL algorithm leveraging the policy-based method (actor) with the value-based approach (critic) to learn the policy and value-function simultaneously (Grondman et al. 2012). In this setup, the actor learns an optimal policy through environment exploration, while the critic evaluates actions to determine long-term rewards (Liu et al. 2021). Utilizing this, Liu et al. (2021) propose FairRec that dynamically maintains accuracy and user fairness in the long term by jointly representing the user preferences (at the item and group level) as the states in an MDP model. They use the Weighted-Proportional Fairness concept, allocating resources (recommendations) proportionally to diverse entities’ demands using a weighting factor. This dual-reward system strategically optimizes both accuracy and fairness in the long term. Separately, Ge et al. (2021) study the long-term fairness of item exposure, focusing on group item labels. They address the challenge of items categorized into groups based on popularity, which dynamically change with user interaction over time, leading to shifting group labels. Using a Demographic Parity-based fairness-constrained policy optimization method, the authors propose an MDP recommendation model based on the actor-critic architecture. However, they introduce an additional critic designed for the demographic parity fairness constraint. On a related note, Morik et al. (2020) explored the rich-get-richer phenomenon in dynamic Learning-to-Rank approaches (Yang and Ai 2021). Their solution, FairCo, ensures meritocratic fairness by allocating exposure to items proportional to their average relevance. FairCo integrates user feedback and dynamically adjusts its ranking based on the fairness constraint, employing a P-Controller approach for modeling, a popular feedback control mechanism (Bequette 2003). In another work, Saito et al. (2020) tackles self-selection bias (Chen et al. 2023), where popular items tend to receive more clicks regardless of actual interest. They propose an ideal loss function optimized for relevance rather than the highest click probability. In this regard, uniform data has also been proven to be an effective approach (Akpinar et al. 2022). This method involves randomly selecting items and ranking them using a uniform distribution. Chen et al. (2023) (Section 4.8) summarizes the various methods that use this approach. Lastly, Ferraro, Serra, and Bauer (2021) investigate gender fairness in collaborative filtering, revealing disparities in exposure between female and male artists. They introduce an iterative re-ranking method that simulates feedback loops and penalizes male artists, resulting in adjusted rankings without significant performance impact.
@@
Discussion
The majority of the long-term fairness evaluations have either been context-specific (e.g., hiring) or in recommendation systems. In recommendation systems, RL methods are favored for modeling dynamics. However, a notable drawback of RL methods is their vulnerability to user tampering (Kasirzadeh and Evans 2023), where recommender systems manipulate user opinions and preferences to boost long-term user engagement through its recommendations. Furthermore, assessing long-term fairness in a multi-stakeholder (Abdollahpouri and Burke 2019) setting characterized by competing interests remains largely unexplored. Some other promising applications include text, where the advent of Large Language Models (LLMs) has facilitated dynamic conversations in various settings like chatbots and healthcare assistants (Gilbert et al. 2023). Studies have indicated that model outputs can be influenced through iterative feedback (Zamfirescu-Pereira et al. 2023; Logan IV et al. 2021), raising fairness concerns. This, coupled with contextual interactions (Zhang et al. 2021), underscores the need to develop evaluation methods capable of assessing long-term fairness to capture these evolving dynamics.
@@ 6 Long-Term Fairness Beyond Dynamic Modeling
Performative prediction refers to the dynamic influence of model decisions on future data distribution. At its core is the concept of a distribution map $D:\Theta\rightarrow\mathcal{P}(X\times Y)$, which captures the dependence of data-generating distribution on the model parameters $\theta\in\Theta$, ignoring other reasons for distribution shift (Hardt and Mendler-Dünner 2023). This abstraction facilitates the modeling of any mechanism for the distribution shift. Miller, Perdomo, and Zrnic (2021) propose a two-stage approach where the learner first estimates a model of the distribution map to capture the unknown parameters of the performative effects and then uses the estimated model to construct the proxy of perturbed performative loss. Next, Mendler-Dünner et al. (2020) categorize the stochastic optimization in performative prediction into greedy deploy and lazy deploy. Greedy deployment involves a sequential program that continually updates model parameters based on current data observations and deploys the updated model. In contrast, lazy deployment progressively approximates repeated risk minimization, collecting multiple rounds of observations to update the model between deployments. The survey by Hardt and Mendler-Dünner (2023) offers a comprehensive review of approaches, which can be extended to encompass long-term fairness settings.
Delayed Impact
@@
Delayed impact represents how the model decisions under fairness interventions may affect the long-term well-being of the groups it aims to protect for (Liu et al. 2018). It can be seen as a case of performative prediction without adopting the formal framework. A couple of works have been brought up to mitigate the delayed impact in the context of multi-armed bandits (Tang, Ho, and Liu 2020) or RL (Weber et al. 2022). The proposed algorithms include two common steps to mitigate the delayed impact. In the first step, the model learns from the historical data through UCB or optimization as a classification problem. In the second step, the delayed impact will be estimated through importance sampling on the realized future data. The estimated delayed impact will be used as a reward signal that can be further used to select the arms or determine the classifier.
Distribution Shift
@@
An alternative research avenue explores dynamic learning environments where data distributions evolve over time. For instance, the profile of loan applicants may shift due to macroeconomic trends or changes in self-selection criteria. This dynamic necessitates continuous fairness maintenance rather than a one-time requirement. A representative work (Almuzaini et al. 2022) proposes a dynamic learning method that proactively corrects the classifier, estimating future data properties and retraining accordingly without direct model-environment interaction. Within this framework, two approaches emerge 1) Adaptive, adapting the classifier to a known target environment, and 2) Robust, training classifiers to ensure fairness under perturbations when no target environment information is available (Barrainkua et al. 2023). This problem setting differs from other works due to the absence of direct environmental feedback. For an in-depth review of works in this area, we refer readers to the survey on fairness under domain shift (Barrainkua et al. 2023).
@@ 7 Open Challenges and Future Works
Evaluating long-term and short-term fairness tension. Evaluating the tension and developing techniques for achieving fairness immediately and maintaining it over an extended period presents a significant challenge. Additionally, while the fairness-utility trade-off has been thoroughly examined in fairness literature (Mehrabi et al. 2021), assessing this long-term trade-off and establishing benchmarks to evaluate different methods remain important directions.
@@
Complexity of real-world dynamics. As we have seen in Sections 2 and 5, long-term fairness necessitates modeling and characterization of dynamics. However, the complexity of real-world scenarios, including limited access to demographic data, presents significant challenges for practitioners to fully understand the interplay of various dynamics and effectively address the discrimination involved. Translating insights from studies focused on specific applications to broader contexts and bridging the gap between empirical fairness evaluations (Gohar, Biswas, and Rajan 2023) and theoretical guarantees (Tang and Zhang 2022) are important and ongoing challenges in the field.
@@
Intersectional identities. Current long-term fairness research has primarily concentrated on single-attribute group identities, such as gender or age. However, numerous studies reveal that bias is a nuanced phenomenon that spans multiple protected groups and intersecting identities (Gohar and Cheng 2023). This complexity poses unique challenges, notably the exponential number of subgroups. Future endeavors should focus on developing methodologies and frameworks that capture and mitigate bias at the intersection of various attributes in long-term settings.
@@
Explainability over time. Explanability in AI has gained tremendous attention due to the need for transparency, which significantly affects acceptance within the broader community (Gross 2007; Shin 2021). Long-term fairness notions, such as effort-based fairness, necessitate using explainable methods. It is vital to empower the impacted population to understand how decisions are made and determine strategies to enhance their qualifications in subsequent instances. Furthermore, incorporating explainability methods in dynamic fairness-aware recommendation environments is another direction (Chen, Zhang, and Qin 2019).
@@
Long-term dynamics in other AI domains. Studies have shown that prompt engineering and fine-tuning in LLMs directly influence the model toward generating the desired outputs (Zamfirescu-Pereira et al. 2023; Logan IV et al. 2021). In real-world scenarios, user interactions are dynamic and may even be strategic, guiding the LLMs in tailoring responses to their needs through repeated feedback. An important research direction is developing evaluation methods that capture these dynamics and assess bias and fairness in LLM (and other large models) outputs. On the other hand, it is also crucial to steer the response of LLMs to consistently align with human values represented by different demographics and cultures (Solaiman et al. 2023) in multi-round dialogues.
@@
Alignment with policy and regulatory framework in the long run. Trustworthy AI has drawn increasing attention from policy and regulatory efforts. For example, the European Commission High-Level Expert Group on AI presented the Ethical Guidelines (European Commission 2019) that put forward a set of 7 key requirements an AI system should meet to be deemed trustworthy. The Organisation for Economic Cooperation and Development also proposed a framework to compare implementation tools for promoting Trustworthy AI (Organisation for Economic Co-operation and Development (2020) OECD). Consequently, it is important to consider long-term fairness together with other aspects of Trustworthy AI.
@@ 8 Conclusion
Ethical Statement
Throughout the process of conducting the project, all authors are aware of and committed to rigorously adhering to AIES Ethics Guidelines. In this work, we review the literature on long-term fairness, examining long-term fairness methods, notions, and evaluations. We also highlight the limitations of current methods and highlight avenues for future research. We hope our work can serve as an informative guide, encouraging researchers to explore the nuances of long-term fairness. We have reviewed the potential ethical implications of our work, and based on our assessment, no foreseeable ethical concerns have been identified.
Appendix A Survey Methodology
We employed several established approaches to conduct a comprehensive search for relevant papers under the umbrella of long-term fairness. Initially, we curated a set of seed papers that have discussed and evaluated long-term fairness problems (Zhang and Liu 2021; Liu et al. 2018; Hu and Chen 2018; D’Amour et al. 2020; Zhang and Liu 2021; Zhang et al. 2020). Following common practice (Desolda et al. 2021; Snyder 2019), we used these papers and domain knowledge to identify and develop a set of keywords to search for relevant work: “long-term fairness,” “dynamic fairness,” “(non-)static fairness,” and “sequential fairness.” Primarily, we searched proceedings of relevant conferences and journals that were likely to publish work that fit the definition of long-term fairness. These included but were not limited to FAccT, NeurIPS, AAAI, ICML, AIES, ICLR, and IJCAI. We also followed citation trails to and from these widely cited papers and querying across Google Scholar and Arxiv to identify additional related work.

### v2

@@ 4.3 Discussion
Compared with the standard static fairness interventions, the summarized mitigation methods mainly contribute to formulating the dynamics of human-model interactions and establishing the corresponding solutions based on the group utility shifts. Long-term fairness under unknown dynamics and/or without leveraging demographic information remains largely unexplored but promising, especially considering privacy laws (Chen et al., 2019a) that limit access to such data. From the perspective of temporal learning, it is positive that long-term fairness is focused on the multi-step sequential learning setting. However, the current research typically only focuses on optimizing the model learner at each time step. Understanding the tensions between ensuring fairness at each time step and cumulatively at the final step needs more investigation. Finally, while RL methods are flexible and learn the dynamics of the environment well, they are data-intensive (Dulac-Arnold et al., 2021; Henderson et al., 2018) and often unsuitable for real-world scenarios. Hence, developing techniques that can efficiently estimate the dynamics is crucial. Notably, the recent advance in performative prediction Perdomo et al. (2020) provides a new perspective on off-policy learning for modeling the interplay between the model policy and agents.
@@ 5 Evaluation of Long-Term Fairness
5.1 Application-Specific Studies
@@
Labor Market. The seminal work of (Hu & Chen, 2018) considers a two-stage dynamic reputational labor market model to capture the long-term and reinforcing nature of disparate outcomes. They propose imposing demographic parity within a temporary labor market to achieve long-term equilibrium in the permanent reputational labor market based on affirmative action. Separately, Celis et al. (2021) investigate the long-term impact of the Rooney Rule (Collins, 2007), a hiring policy aimed at mitigating implicit bias by mandating decision-makers to include at least one candidate from an underrepresented group. They employ a mathematical model of implicit bias, featuring two candidate groups, one of which is underrepresented. Each candidate possesses both a true, latent utility and an observed utility, subject to bias by a multiplicative factor. Candidates are then shortlisted by a selection panel based on observed utility. Their findings demonstrate that the rule progressively enhances minority hiring while maintaining overall utility.
Predictive Policing Ensign et al. (2018) explore resource allocation in predictive policing (Mohler et al., 2015), which deploys patrol officers based on historical crime data. They identify a feedback phenomenon present in similar algorithms across contexts like recidivism prediction and lending, where prediction outcomes heavily influence algorithm feedback. Using the Pólya urn model, they explain this feedback occurrence and propose black-box solutions, such as reweighing input incidents. Elzayn et al. (2019), go one step further by proposing an algorithm which, under simplification assumptions, converges to an optimally fair resource allocation, even in the absence of known stationary ground-truth crime rates, avoiding the runaway feedback loop effect.
5.2 Long-Term Fairness in Recommender Systems
Long-term fairness considerations naturally align with Recommender Systems due to evolving user interactions and item popularity, impacting system performance Liu et al. (2021). Consequently, the RL framework is favored for its ability to model interactive dynamics. The works mostly focus on 1) user-side fairness, addressing equitable treatment for users, and 2) item-side fairness, ensuring unbiased item recommendation stemming from popularity bias due to disproportionate exposure Chen et al. (2023).
@@ 5.2 Long-Term Fairness in Recommender Systems
Long-term fairness efforts in recommendation systems primarily target item-side fairness, addressing the challenge by modeling the sequential interactions between consumers and recommender systems using an MDP framework. A few works in this direction adopt an Actor-Critic RL algorithm leveraging the policy-based method (actor) with the value-based approach (critic) to learn the policy and value-function simultaneously Grondman et al. (2012). In this setup, the actor learns an optimal policy through environment exploration, while the critic evaluates actions to determine long-term rewards Liu et al. (2021). Utilizing this, Liu et al. (2021) propose FairRec that dynamically maintains accuracy and user fairness in the long term by jointly representing the user preferences (at the item and group level) as the states in an MDP model. They use the Weighted-Proportional Fairness concept, allocating resources (recommendations) proportionally to diverse entities’ demands using a weighting factor. This dual-reward system strategically optimizes both accuracy and fairness in the long term. Separately, Ge et al. (2021) study the long-term fairness of item exposure, focusing on group item labels. They address the challenge of items categorized into groups based on popularity, which dynamically change with user interaction over time, leading to shifting group labels. Using a Demographic Parity-based fairness-constrained policy optimization method, the authors propose an MDP recommendation model based on the actor-critic architecture. However, they introduce an additional critic designed for the demographic parity fairness constraint. On a related note, Morik et al. (2020) explored the rich-get-richer phenomenon in dynamic Learning-to-Rank approaches Yang & Ai (2021). Their solution, FairCo, ensures meritocratic fairness by allocating exposure to items proportional to their average relevance. FairCo integrates user feedback and dynamically adjusts its ranking based on the fairness constraint, employing a P-Controller approach for modeling, a popular feedback control mechanism (Bequette, 2003). In another work, Saito et al. (2020) tackles self-selection bias (Chen et al., 2023), where popular items tend to receive more clicks regardless of actual interest. They propose an ideal loss function optimized for relevance rather than the highest click probability. In this regard, uniform data has also been proven to be an effective approach Akpinar et al. (2022). This method involves randomly selecting items and ranking them using a uniform distribution. Chen et al. (2023) (Section 4.8) summarizes the various methods that use this approach. Lastly, Ferraro et al. (2021) investigate gender fairness in collaborative filtering, revealing disparities in exposure between female and male artists. They introduce an iterative re-ranking method that simulates feedback loops and penalizes male artists, resulting in adjusted rankings without significant performance impact.
@@
5.3 Discussion
The majority of the long-term fairness evaluations have either been context-specific (e.g., hiring) or in recommendation systems. In recommendation systems, RL methods are favored for modeling dynamics. However, a notable drawback of RL methods is their vulnerability to user tampering (Kasirzadeh & Evans, 2023), where recommender systems manipulate user opinions and preferences to boost long-term user engagement through its recommendations. Furthermore, assessing long-term fairness in a multi-stakeholder (Abdollahpouri & Burke, 2019) setting characterized by competing interests remains largely unexplored. Some other promising applications include text, where the advent of Large Language Models (LLMs) has facilitated dynamic conversations in various settings like chatbots and healthcare assistants (Gilbert et al., 2023). Studies have indicated that model outputs can be influenced through iterative feedback (Zamfirescu-Pereira et al., 2023; Logan IV et al., 2021), raising fairness concerns. This, coupled with contextual interactions (Zhang et al., 2021), underscores the need to develop evaluation methods capable of assessing long-term fairness to capture these evolving dynamics.
@@ 6.1 Performative Prediction
Performative prediction refers to the dynamic influence of model decisions on future data distribution. At its core is the concept of a distribution map $D:\Theta\rightarrow\mathcal{P}(X\times Y)$, which captures the dependence of data-generating distribution on the model parameters $\theta\in\Theta$, ignoring other reasons for distribution shift (Hardt & Mendler-Dünner, 2023). This abstraction facilitates the modeling of any mechanism for the distribution shift. Miller et al. (2021) propose a two-stage approach where the learner first estimates a model of the distribution map to capture the unknown parameters of the performative effects and then uses the estimated model to construct the proxy of perturbed performative loss. Next, Mendler-Dünner et al. (2020) categorize the stochastic optimization in performative prediction into greedy deploy and lazy deploy. Greedy deployment involves a sequential program that continually updates model parameters based on current data observations and deploys the updated model. In contrast, lazy deployment progressively approximates repeated risk minimization, collecting multiple rounds of observations to update the model between deployments. The survey by Hardt & Mendler-Dünner (2023) offers a comprehensive review of approaches, which can be extended to encompass long-term fairness settings.
6.2 Delayed Impact
@@
Delayed impact represents how the model decisions under fairness interventions may affect the long-term well-being of the groups it aims to protect for Liu et al. (2018). It can be seen as a case of performative prediction without adopting the formal framework. A couple of works have been brought up to mitigate the delayed impact in the context of multi-armed bandits Tang et al. (2020) or RL Weber et al. (2022). The proposed algorithms include two common steps to mitigate the delayed impact. In the first step, the model learns from the historical data through UCB or optimization as a classification problem. In the second step, the delayed impact will be estimated through importance sampling on the realized future data. The estimated delayed impact will be used as a reward signal that can be further used to select the arms or determine the classifier.
6.3 Distribution Shift
@@
An alternative research avenue explores dynamic learning environments where data distributions evolve over time. For instance, the profile of loan applicants may shift due to macroeconomic trends or changes in self-selection criteria. This dynamic necessitates continuous fairness maintenance rather than a one-time requirement. A representative work Almuzaini et al. (2022) proposes a dynamic learning method that proactively corrects the classifier, estimating future data properties and retraining accordingly without direct model-environment interaction. Within this framework, two approaches emerge 1) Adaptive, adapting the classifier to a known target environment, and 2) Robust, training classifiers to ensure fairness under perturbations when no target environment information is available (Barrainkua et al., 2023). This problem setting differs from other works due to the absence of direct environmental feedback. For an in-depth review of works in this area, we refer readers to the survey on fairness under domain shift Barrainkua et al. (2023).
@@ 7 Open Challenges and Future Works
Evaluating long-term and short-term fairness tension. Evaluating the tension and developing techniques for achieving fairness immediately and maintaining it over an extended period presents a significant challenge. Additionally, while the fairness-utility trade-off has been thoroughly examined in fairness literature (Mehrabi et al., 2021), assessing this long-term trade-off and establishing benchmarks to evaluate different methods remain important directions.
@@
Complexity of real-world dynamics. As we have seen in Sections 2.3 and 5, long-term fairness necessitates modeling and characterization of dynamics. However, the complexity of real-world scenarios, including limited access to demographic data, presents significant challenges for practitioners to fully understand the interplay of various dynamics and effectively address the discrimination involved. Translating insights from studies focused on specific applications to broader contexts and bridging the gap between empirical fairness evaluations (Gohar et al., 2023) and theoretical guarantees (Tang & Zhang, 2022) are important and ongoing challenges in the field.
@@
Intersectional identities. Current long-term fairness research has primarily concentrated on single-attribute group identities, such as gender or age. However, numerous studies reveal that bias is a nuanced phenomenon that spans multiple protected groups and intersecting identities Gohar & Cheng (2023). This complexity poses unique challenges, notably the exponential number of subgroups. Future endeavors should focus on developing methodologies and frameworks that capture and mitigate bias at the intersection of various attributes in long-term settings.
@@
Explainability over time. Explanability in AI has gained tremendous attention due to the need for transparency, which significantly affects acceptance within the broader community Gross (2007); Shin (2021). Long-term fairness notions, such as effort-based fairness, necessitate using explainable methods. It is vital to empower the impacted population to understand how decisions are made and determine strategies to enhance their qualifications in subsequent instances. Furthermore, incorporating explainability methods in dynamic fairness-aware recommendation environments is another direction Chen et al. (2019b).
@@
Long-term dynamics in other AI domains. Studies have shown that prompt engineering and fine-tuning in LLMs directly influence the model toward generating the desired outputs Zamfirescu-Pereira et al. (2023); Logan IV et al. (2021). In real-world scenarios, user interactions are dynamic and may even be strategic, guiding the LLMs in tailoring responses to their needs through repeated feedback. An important research direction is developing evaluation methods that capture these dynamics and assess bias and fairness in LLM (and other large models) outputs. On the other hand, it is also crucial to steer the response of LLMs to consistently align with human values represented by different demographics and cultures (Solaiman et al., 2023) in multi-round dialogues.
@@
Alignment with policy and regulatory framework in the long run. Trustworthy AI has drawn increasing attention from policy and regulatory efforts. For example, the European Commission High-Level Expert Group on AI presented the Ethical Guidelines (European Commission, 2019) that put forward a set of 7 key requirements an AI system should meet to be deemed trustworthy. The Organisation for Economic Cooperation and Development also proposed a framework to compare implementation tools for promoting Trustworthy AI (Organisation for Economic Co-operation and Development (2020), OECD). Consequently, it is important to consider long-term fairness together with other aspects of Trustworthy AI.
@@ 8 Conclusion
Appendix A Appendix
A.1 Survey Methodology
We employed several established approaches to conduct a comprehensive search for relevant papers under the umbrella of long-term fairness. Initially, we curated a set of seed papers that have discussed and evaluated long-term fairness problems (Zhang & Liu, 2021; Liu et al., 2018; Hu & Chen, 2018; D’Amour et al., 2020; Zhang & Liu, 2021; Zhang et al., 2020). Following common practice (Desolda et al., 2021; Snyder, 2019), we used these papers and domain knowledge to identify and develop a set of keywords to search for relevant work: “long-term fairness," “dynamic fairness," “(non-)static fairness," and “sequential fairness." Primarily, we sear‘ched proceedings of relevant conferences and journals that were likely to publish work that fit the definition of long-term fairness. These included but were not limited to FAccT, NeurIPS, AAAI, ICML, AIES, ICLR, and IJCAI. We also followed citation trails to and from these widely cited papers and querying across Google Scholar and Arxiv to identify additional related work.